OpenStudy (anonymous):

2 slope problems PLEASE help Write the equation of the line with a slope of -1/3 and passing through the point (6, -4). Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (anonymous):

For the first equation, use the point slope formula and plug in the given data

OpenStudy (anonymous):

y+4=(-1/3)(x-6) ----->Now rearrange this if it is required. You can leave it like this since it is correct.

OpenStudy (anonymous):

And for the second part, find the slope between the two points: (-6,1) and (-4,2).. And use one of the points to form the equation.

OpenStudy (anonymous):

it has 2 b completely simplified and 4 the first 1 got y=-1/3x-6

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

It would have helped if you would spell out the words. And no, it is not y=-1/3x-6. Instead it is y=-1/3x-2

OpenStudy (anonymous):

oh ok I see what I did wrong and sry im trying 2 type with 1 hand and it takes longer so im using short cuts

OpenStudy (anonymous):

For the second part, can you find the slope between the two points?

OpenStudy (anonymous):

The slope would be 1/2. Now pick either point and write the equation.

OpenStudy (anonymous):

y-2=1/2(x+4)----> Now rearrange this and get y=1/2x+4

OpenStudy (anonymous):

Hope this helped.

OpenStudy (anonymous):

ok thanks :) do you understand this problem : Write the equation of the line with an undefined slope, passing through the point (2, 5).

OpenStudy (anonymous):

I don't understand how u can write an undefined slope

OpenStudy (anonymous):

undefined slope means the line is vertical. Therefore, the equation for the line would be y=2

OpenStudy (anonymous):

wait, x=2.

OpenStudy (anonymous):

so you don't do anything with the 5?

OpenStudy (anonymous):

y=2 would be horizontal, so it would be x=2

OpenStudy (anonymous):

5 is a point above the 2. Only the first variable matters, which is 2.

OpenStudy (anonymous):

oh ok :) thank u sooo much

OpenStudy (anonymous):